OpenStudy (anonymous):

Find x + 50 if x = 50 and y = 2 y

OpenStudy (anonymous):

50 52 75 78

Vocaloid (vocaloid):

can you try typing out the problem again? it got a bit scrambled

OpenStudy (anonymous):

x/y +50 if x = 50 and y = 2

OpenStudy (anonymous):

Question: Wendy is looking over some data regarding the strength, measured in pascals ( pa ) , of some rope and how the strength relates to the number of woven strands in the rope. The data represented by the exponential function f ( x ) = 2x , where x is the number of woven strands. Explain how she can convert this equation to logarithmic function when strength is 256 pascals

OpenStudy (anonymous):

its 52

OpenStudy (anonymous):

thanks guys

OpenStudy (texaschic101):

x/y + 50....when x = 50 and y = 2 just a matter of subbing... 50/2 + 50 = 25 + 50 = 75